# Quasi-boolean groups

### v2

@@ 1. Introduction
Following Haran and Jarden [5] we say that a profinite group $G$ is real projective if $G$ has an open subgroup without $2$-torsion, and if every real embedding problem for $G$ has a solution.
By the work of Haran–Jarden [5], real projective groups play an important role in Galois theory as they are the absolute Galois groups of pseudo real closed fields. In fact, it follows from the work of Haran [4] and Haran–Jarden [5] that the absolute Galois groups of fields with virtual cohomological dimension at most one, which is a slightly larger class of fields, are real projective (see [8, Corollary 2.4]). For a real projective group $G$, we show in [8, Theorem 1.3] that the differential graded algebra $C^{\bullet}(G,\mathbb{F}_{2})$ of continuous cochains is formal, i.e., $C^{\bullet}(G,\mathbb{F}_{2})$ is quasi-isomorphic as differential graded algebras to its cohomology algebra. Roughly speaking, this means that the cohomology algebra of a real projective group already contains all the information of the differential graded algebra $C^{\bullet}(G,\mathbb{F}_{2})$. The purpose of the present paper is to show that, in fact, the maximal pro-$2$ quotient of a real projective group can be reconstructed entirely from the mod $2$ cohomology ring. In particular, we show that pro-$2$ real projective groups can be reconstructed entirely from their mod $2$ cohomology ring. In order to further describe our results we recall the following terminology from [8].
@@
We now describe our main results in more detail. To do so we need the following constructions and terminology:
@@
The free product $G_{1}*_{p}G_{2}$ of two pro-$p$ groups $G_{1},G_{2}$ is the following: let $G_{1}*G_{2}$ be the discrete free product of $G_{1}$ and $G_{2}$ and let $\mathcal{N}$ be the family of normal subgroups $N$ of $G$ such that $(G_{1}*G_{2})/N$ is a finite $p$-group and $N\cap G_{1},N\cap G_{2}$ are open subgroups of $G_{1},G_{2}$ respectively. Then
@@
Next we are going to define, following [2], free pro-$2$ products of order two groups over topological spaces. For every topological space $X$ let $\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ denote the group which is freely generated by the elements of $X$, subject to the relation that these elements are involutions, and let $\mathcal{N}$ be the family of normal subgroups $N$ of $\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ such that $\bigast_{X}\mathbb{Z}/2\mathbb{Z}/N$ is a finite $2$-group and the composition of the natural inclusion $\iota\colon X\to\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ and the quotient homomorphism $\bigast_{X}\mathbb{Z}/2\mathbb{Z}\to\bigast_{X}\mathbb{Z}/2\mathbb{Z}/N$ is continuous with respect to the discrete topology on $\bigast_{X}\mathbb{Z}/2\mathbb{Z}/N$. Then we set
@@
We say that a pro-$2$ group is a Boolean group if it is isomorphic to $\mathbb{B}(X)$ for some topological space $X$. (We will see in Proposition 6.5 below that we may assume without the loss of generality that $X$ is profinite.) We say that a pro-$2$ group is a quasi-Boolean group if it is the free product of a free pro-$2$ group (in the sense of [12, 1.5 on pages 7–8]) and a Boolean group. We say that a pro-$2$ group is a cohomologically Boolean group if its mod $2$ cohomology is a graded Boolean algebra. We say that a pro-$2$ group is a cohomologically quasi-Boolean group if its mod $2$ cohomology is the connected sum of a dual algebra and a graded Boolean algebra.
We can now state our main results:
@@ 4. $(iv)$
Perhaps the most interesting feature of this theorem is, compared to the results of the paper [6], that it incorporates a purely cohomological characterisation of these pro-$2$ groups. Our proof is a bit more involved than it might be anticipated; it uses results of Haran–Jarden in the arithmetic of fields, for example a group-theoretical characterisation of real projective groups, a theorem on the existence of sections of profinite principle $G$-bundles, and profinite versions of two theorems of Quillen on group cohomology. We will also show the following
@@ 2. $(ii)$
Using our results we also demonstrate that quasi-Boolean groups can be reconstructed from their mod $2$ cohomology. For every set $Y$, let $F(Y)$ denote the free pro-$2$ group as defined in [12, Section 1.5 on pages 7–8]. Let $G$ be a quasi-Boolean group such that $H^{\bullet}(G,\mathbb{F}_{2})$ is the connected sum of a dual algebra $D^{\bullet}$ and a graded Boolean algebra $B^{\bullet}$ associated to the Boolean ring $B$. Let $Y$ be a basis of $D^{1}$ and let $X$ be the spectrum of $B$.
@@
The pro-$2$ group $G$ is isomorphic to $F(Y)*_{2}\mathbb{B}(X)$.
@@
In Section 2 we give a modern exposition of the theory of Boolean rings, including Stone duality, using now standard tools from commutative algebra. We cover some background material on profinite spaces, including the profinite completion functor, in the Section 3. In Section 4 we prove that profinite principal $G$-bundles have sections, a result originally announced by Morel in [7] without proof. We prove that the class of pro-$2$ real projective groups and the maximal pro-$2$ quotients of real projective groups are the same in Section 5, the key step being a simple group-theoretical lemma. In Section 6 we show that the class of quasi-Boolean and pro-$2$ real projective groups are the same, heavily relying on the main results of the previous sections and several theorems of the paper [6]. Then we present profinite versions of some classical results of Quillen on the cohomology of groups, partially following the suggestion at the end of [11], and use these results to derive a local-global principle for the cohomology of cohomologically quasi-Boolean groups analogous to Scheiderer’s theorem for real projective groups in Section 7. In section 8 we prove the main results using cohomological obstruction theory for central embedding problems and the local-global principle of the previous section.
@@ Acknowledgement
This paper was written as a response to one of the questions asked by the referee to our paper [8]. We are very thankful for the anonymous referee’s comments. Both authors gratefully acknowledge the partial support by the RCN Project No. 313472 Equations in Motivic Homotopy.
@@ 2. A primer on Boolean rings
In this section we recall and prove the results we need on Boolean rings. In particular, we deduce Stone duality in Theorem 2.10.
@@
Examples 2.2.
The field with two elements $\mathbb{F}_{2}$ is a Boolean ring. In fact, since $x(1-x)=0$ for all $x$ in a Boolean ring, $\mathbb{F}_{2}$ is the only Boolean integral domain. The direct product of Boolean rings is Boolean, and so, for every set $X$, the direct product ring:
@@ 2. $(ii)$
    every prime ideal $\mathfrak{p}$ is maximal, and $R/\mathfrak{p}$ is the field with two elements,
@@ 4. $(iv)$
Since
$2x=(2x)^{2}=4x^{2}=4x,$
we get that $02x=$ 0 by subtracting $2x$ from both sides. Now let $\mathfrak{p}$ be a prime ideal in $R$. Then the quotient $R/\mathfrak{p}$ is a Boolean ring. For every $x\in R/\mathfrak{p}$, we have $x(1-x)=0$ which implies that $x=0$ or $x=1$ since $R/\mathfrak{p}$ is an integral domain. Claim $(ii)$ follows. Note that
@@
Hence $x,y\in(x+y-xy)$. Since $x+y-xy\in(x,y)$, claim $(iii)$ is clear. Let $I=(x_{1},x_{2},\ldots,x_{n})$ be an finitely generated ideal of $R$. Since
@@
we may assume by induction on $n$ that $I=(x,y)$ for some $x,y\in R$. The claim now follows from part $(iii)$. ∎
@@
The spectrum $\mathrm{Spec}(R)$ of a Boolean ring is compact and totally separated.
@@
Since the spectrum of a commutative ring with a unity is compact, the same holds for $\mathrm{Spec}(R)$, too. Recall that a topological space $X$ is totally separated if for any two distinct points $x,y\in X$ there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$. Now let $\mathfrak{p},\mathfrak{q}\in\mathrm{Spec}(R)$ be two distinct points. Since they are maximal ideals, there is an $x\in R$ such that $x\in\mathfrak{p}$ and $x\not\in\mathfrak{q}$. Since $R/\mathfrak{p}=\mathbb{F}_{2}$ by part $(ii)$ of Proposition 2.3, the former is equivalent to $1-x\not\in\mathfrak{p}$. As usual for every $f\in R$ let $D(f)\subseteq\mathrm{Spec}(R)$ denote the open subset
@@
Notation 2.5.
Let $R$ be a Boolean ring. Then for every $a\in R$ the corresponding section of the structure sheaf of $\mathrm{Spec}(R)$ is a continuous function
@@
by part $(ii)$ of Proposition 2.3, and the furnished map
@@
For every $x\in R$ we have $x^{n}=x$ by induction, so if $x$ is nilpotent, then it is zero. Therefore the nilradical of $R$ is zero, so by Krull’s theorem $\sigma$ is injective. So we only need to show that $\sigma$ is surjective. Let $f\colon\mathrm{Spec}(R)\to\mathbb{F}_{2}$ be a continuous function. Then the set
$D(f)=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid f(x)=1\}$
is a closed subset, so it is compact. But it is also open, so it can be covered by open subsets of the form $D(a)$, where $a\in R$ by the definition of the topology of the spectrum of rings. Since $D(f)$ is compact, it can be covered by finitely many such, so
@@
$D(a)$ =$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a\not\in\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a)\not\subseteq\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a_{1},\ldots,a_{n})\not\subseteq\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a_{i}\not\in\mathfrak{p}\textrm{ for some $i$}\}$
@@
so $D(f)=D(a)$. Since $f$ and $a$ take values in $\mathbb{F}_{2}$, we get that $f=a$. ∎
Notation 2.7.
@@
For every $x\in\mathbf{B}(X)$ the pre-image
$\beta^{-1}(D(x))=\{p\in X\mid x(p)=1\}$
is open. Since the sets $\{D(x)\mid x\in\mathbf{B}(X)\}$ form a sub-basis of $\mathrm{Spec}(\mathbf{B}(X))$, we get that $\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X))$ is continuous.
@@
Since $X$ is compact and $\mathrm{Spec}(\mathbf{B}(X))$ is Hausdorff, it will be sufficient to show that $\beta$ is a bijection. Let $x,y\in X$ be two distinct points. Since $X$ is totally separated, so there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$. Let $f\colon X\to\mathbb{F}_{2}$ be the characteristic function of $U$. It is in $\mathcal{B}(X)$ since the complement of $U$ is open, too. Clearly $f\in\beta(y)$, but $f\not\in\beta(x)$, so $\beta$ is injective.
For every ideal $I\triangleleft\mathbf{B}(X)$ let $Z(I)\subseteq X$ denote the closed subset
@@
so if this set is empty, then $f$ is the identically one function, and hence $I=(1)=\mathcal{B}(X)$ is not proper, a contradiction. Next consider the case when $I$ is finitely generated; then it is principal by part $(iv)$ of Proposition 2.3, so $Z(I)$ is non-empty by the above. Finally, consider the general case. Then $Z(I)$ is the intersection of sets of the form $Z(J)$ where $J$ is a finitely generated ideal of $I$. Since the latter collection of sets is closed under finite intersections, and each member is non-empty by the above, we get $Z(I)$ is also non-empty, since $X$ is compact.
Now let $\mathfrak{m}\triangleleft\mathbf{B}(X)$ be a maximal ideal. By the above there is an $x\in Z(\mathfrak{m})$. Clearly $\beta(x)\supseteq\mathfrak{m}$, but $\mathfrak{m}$ is maximal, and hence $\beta(x)=\mathfrak{m}$. Therefore $\beta$ is surjective, too. ∎
Notation 2.9.
@@
which is well-defined as we saw in Examples 2.2, and
@@
Let $R$ be a finite Boolean ring. Then the following are equivalent:

### v3

@@ Abstract.
Both authors were partially supported by RCN Project No. 313472 Equations in Motivic Homotopy, and the project Pure Mathematics in Norway funded by the Trond Mohn Foundation.
@@ 1. Introduction
Following Haran and Jarden [9, page 450] we say that a profinite group $G$ is real projective if the subset $\mathrm{Inv}(G)$ of involutions is closed in $G$ and if every real embedding problem for $G$ has a solution.
Remark 1.2.
By [9, Remark 7.6 and Proposition 7.7 on page 473], the condition that $\mathrm{Inv}(G)$ is closed in $G$ is equivalent to the condition that $G$ has an open subgroup without $2$-torsion which is used in [15, Definition 1.1]. To see that these conditions are equivlaent, let $g\in G$ be an element which belongs to the closure of $\mathrm{Inv}(G)$ satisfies $g^{2}=1$. Then $g\in\mathrm{Inv}(G)$ or $g=1$. Thus, $\mathrm{Inv}(G)$ is closed in $G$ if and only if $G$ has an open subgroup $U$ such that $U\cap\mathrm{Inv}(G)=\emptyset$.
By the work of Haran–Jarden [9], real projective groups play an important role in Galois theory as they are the absolute Galois groups of pseudo real closed fields. In fact, it follows from the work of Haran [8] and Haran–Jarden [9] that the absolute Galois groups of fields with virtual cohomological dimension at most one, which is a slightly larger class of fields, are real projective (see [15, Corollary 2.4]). For a real projective group $G$, we show in [15, Theorem 1.3] that the differential graded algebra $C^{\bullet}(G,\mathbb{F}_{2})$ of continuous cochains is formal, i.e., $C^{\bullet}(G,\mathbb{F}_{2})$ is quasi-isomorphic as differential graded algebras to its cohomology algebra. Roughly speaking, this means that the cohomology algebra of a real projective group already contains all the information of the differential graded algebra $C^{\bullet}(G,\mathbb{F}_{2})$. The purpose of the present paper is to show that, in fact, the maximal pro-$2$ quotient of a real projective group can be reconstructed entirely from the mod $2$ cohomology ring. In particular, we show that pro-$2$ real projective groups can be reconstructed entirely from their mod $2$ cohomology ring. In order to further describe our results we recall the following terminology from [15].
@@
We now describe our main results in more detail. To do so we need the following constructions and terminology. First we recall free products.
@@
Let $G_{1}$ and $G_{2}$ be two pro-$p$ groups. Let $G_{1}*G_{2}$ be the discrete free product of $G_{1}$ and $G_{2}$ and let $\mathcal{N}$ be the family of normal subgroups $N$ of $G$ such that $(G_{1}*G_{2})/N$ is a finite $p$-group and $N\cap G_{1},N\cap G_{2}$ are open subgroups of $G_{1},G_{2}$ respectively. Then
@@
is called the free pro-$p$ product of $G_{1}$ and $G_{2}$.
We note that the free pro-$p$ product of $G_{1}$ and $G_{2}$ satisfies the usual universal property with respect to continuous homomorphisms from $G_{1}$ and $G_{2}$ to pro-$p$ groups (see e.g. [11, Proposition 3.1.1]). Next we are going to define free products over topological spaces where we refer to [3], [6], [7], [8], [10], [11, Chapter 4], [12], [17, Chapter 5] for properties and variations of free products of profinite groups.
@@
For every topological space $X$, let $\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ denote the group which is freely generated by the elements of $X$, subject to the relation that these elements are involutions, and let $\mathcal{N}$ be the family of normal subgroups $N$ of $\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ such that $(\bigast_{X}\mathbb{Z}/2\mathbb{Z})/N$ is a finite $2$-group and the composition of the natural inclusion $\iota\colon X\to\bigast_{X}\mathbb{Z}/2\mathbb{Z}$ and the quotient homomorphism $\bigast_{X}\mathbb{Z}/2\mathbb{Z}\to(\bigast_{X}\mathbb{Z}/2\mathbb{Z})/N$ is continuous with respect to the discrete topology on $(\bigast_{X}\mathbb{Z}/2\mathbb{Z})/N$. Then we set
@@
and refer to $\mathbb{B}(X)$ as the free pro-$2$ product of order two groups over $X$.
Finally, we are going to use the following terminology for pro-$2$ groups, where we refer to [21, 1.5 on pages 7–8] for the notion of a free pro-$2$ group.
@@
We say that a pro-$2$ group is a Boolean group if it is isomorphic to $\mathbb{B}(X)$ for some topological space $X$. We say that a pro-$2$ group is a quasi-Boolean group if it is the free product of a free pro-$2$ group and a Boolean group. We say that a pro-$2$ group is a cohomologically Boolean group if its mod $2$ cohomology is a graded Boolean algebra. We say that a pro-$2$ group is a cohomologically quasi-Boolean group if its mod $2$ cohomology is the connected sum of a dual algebra and a graded Boolean algebra.
For a Boolean group $\mathbb{B}(X)$, we will see in Proposition 6.5 that we may assume without the loss of generality that $X$ is profinite. We note that a Boolean group is the free product over a constant sheaf of a group of order $2$ and a profinite space $X$ in the sense of [11, Chapter 4], and quasi-Boolean groups are the real free groups in the category of pro-$2$ groups in the sense of [7] and [10]. We can now state our main results.
@@ 4. $(iv)$
We note that the equivalence of $(ii)$ and $(iii)$ follows from the arguments developed in [10, Section 3], and the equivalence of $(i)$ and $(ii)$ has already been proven in [10, Corollary 3.3] and [7, Proposition 4.2] (assuming that $X$ is a closed system of representatives of the conjugacy classes of involutions in $G$). We give an alternative proof of the equivalence of $(i)$ and $(ii)$ with no assumption on $X$ using results of [10] since this equivalence will follow from results we need to develop for the remaining equivalence with $(iv)$. Hence the most interesting feature of Theorem 1.7 beyond the results of [7] and [10] is that it incorporates a purely cohomological characterisation of these pro-$2$ groups. Our proof is a bit more involved than it might be anticipated. The proof uses results of Haran–Jarden in the arithmetic of fields, for example a group-theoretical characterisation of real projective groups, a theorem on the existence of sections of profinite principle $G$-bundles, and profinite versions of two theorems of Quillen on group cohomology. We will also show the following:
@@ 2. $(ii)$
Using our results we then demonstrate that quasi-Boolean groups can, in fact, be reconstructed from their mod $2$ cohomology. For every set $Y$, let $F(Y)$ denote the free pro-$2$ group as defined in [21, Section 1.5 on pages 7–8].
@@
Let $G$ be a quasi-Boolean pro-$2$ group such that $H^{\bullet}(G,\mathbb{F}_{2})$ is the connected sum of a dual algebra $D^{\bullet}$ and a graded Boolean algebra $B^{\bullet}$ associated to the Boolean ring $B$. Let $Y$ be a basis of $D^{1}$ and let $X$ be the spectrum of $B$. Then $G$ is isomorphic to $F(Y)*_{2}\mathbb{B}(X)$.
@@
In Section 2 we give a modern exposition of the theory of Boolean rings, including Stone duality, using now standard tools from commutative algebra. We cover some background material on profinite spaces, including the profinite completion functor, in the Section 3. In Section 4 we prove that profinite principal $G$-bundles have sections, a result originally announced by Morel in [13] without proof. We prove that the class of pro-$2$ real projective groups and the maximal pro-$2$ quotients of real projective groups are the same in Section 5, the key step being a simple group-theoretical lemma. In Section 6 we show that the class of quasi-Boolean and pro-$2$ real projective groups are the same, heavily relying on the main results of the previous sections and several theorems of [10]. Then, in Section 7, we present profinite versions of some classical results of Quillen on the cohomology of groups, partially following the suggestion at the end of [20], and use these results to derive a local-global principle for the cohomology of cohomologically quasi-Boolean groups analogous to Scheiderer’s theorem for real projective groups. In Section 8 we prove the main results using cohomological obstruction theory for central embedding problems and the local-global principle of the previous section.
@@ Acknowledgement
This paper was written as a response to one of the questions asked by the referee to our paper [15]. We are very thankful for the anonymous referee’s comments. We thank the anonymous referee of the present paper for many comments and suggestions that helped to improve the manuscript.
@@ 2. A primer on Boolean rings
In this section we recall and prove the results we need on Boolean rings. In particular, we deduce Stone duality in Theorem 2.11. We note that the content of this section is very similar to the content of [15, Appendix A]. We include it here for convenience of the reader.
@@
Example 2.2.
The field with two elements $\mathbb{F}_{2}$ is a Boolean ring. In fact, since x(1x)=0x(1-x)=0 for all $x$ in a Boolean ring, $\mathbb{F}_{2}$ is the only Boolean integral domain. The direct product of Boolean rings is Boolean, and so, for every set $X$, the direct product ring:
@@ 2. $(ii)$
    every prime ideal $\mathfrak{p}$ is maximal, and $R/\mathfrak{p}$ is the field with two elements;
@@ 4. $(iv)$
Since $2x=(2x)^{2}=4x^{2}=4x$, we get that $2x=0$ by subtracting $2x$ from both sides. Now let $\mathfrak{p}$ be a prime ideal in $R$. Then the quotient $R/\mathfrak{p}$ is a Boolean ring. For every $x\in R/\mathfrak{p}$, we have x(1x)=0x(1-x)=0 which implies that $x=0$ or $x=1$ since $R/\mathfrak{p}$ is an integral domain. Claim $(ii)$ follows. Note that
@@
Hence $x,y\in(x+y-xy)$. Since $x+y-xy\in(x,y)$, claim $(iii)$ is clear. Let $I=(x_{1},x_{2},\ldots,x_{n})$ be a finitely generated ideal of $R$. Since
@@
we may assume by induction on $n$ that $I=(x,y)$ for some $x,y\in R$. Claim $(iv)$ now follows from part $(iii)$. ∎
Definition 2.4.
A topological space $X$ is called totally separated if for all $x\neq y$ in $X$ there exists a closed and open set $U\subset X$ such that $x\in U$ and $y\notin U$. Equivalently, $X$ is totally separated if for any two distinct points $x,y\in X$ there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$.
@@
The spectrum $\mathrm{Spec}(R)$ of a Boolean ring with the Zariski topology is compact and totally separated.
@@
By [5, Proposition I.1.1.10], the spectrum of a commutative ring with a unity with the Zariski topology is compact. Thus, $\mathrm{Spec}(R)$ is compact. Now let $\mathfrak{p},\mathfrak{q}\in\mathrm{Spec}(R)$ be two distinct points. Since they are maximal ideals by part $(ii)$ of Proposition 2.3, there is an $x\in R$ such that $x\in\mathfrak{p}$ and $x\not\in\mathfrak{q}$. Since $R/\mathfrak{p}=\mathbb{F}_{2}$ by part $(ii)$ of Proposition 2.3, the former is equivalent to $1-x\not\in\mathfrak{p}$. As usual, for every $f\in R$, let $D(f)\subseteq\mathrm{Spec}(R)$ denote the open subset
@@
Notation 2.6.
Let $R$ be a Boolean ring. Then for every $a\in R$ the corresponding section of the structure sheaf of $\mathrm{Spec}(R)$ is the function
@@
defined by $\sigma(a)(\mathfrak{p})=0$ if $a+\mathfrak{p}=\mathfrak{p}$, i.e., $a\in\mathfrak{p}$, and $\sigma(a)(\mathfrak{p})=1$ if $a\notin\mathfrak{p}$. For every $a\in R$, $\sigma(a)$ is continuous since $\{\mathfrak{p}\in\mathrm{Spec}(R)~|~\sigma(a)(\mathfrak{p})=1\}=D(a)$ and $\{\mathfrak{p}\in\mathrm{Spec}(R)~|~\sigma(a)(\mathfrak{p})=0\}=D(1-a)$ are open, because $a\in\mathfrak{p}$ if and only if $1-a\notin\mathfrak{p}$, by part $(ii)$ of Proposition 2.3. The induced map
@@
For every $x\in R$, we have $x^{n}=x$ by induction, so if $x$ is nilpotent, then it is zero. Therefore the nilradical of $R$ is zero. By [1, Proposition 1.8 on page 5], this shows that the intersection of all prime ideals in $R$ is zero. Since $\sigma(a)$ is the zero map if and only if $a$ lies in the intersection of all prime ideals by definition, this shows that $\sigma$ is injective. It remains to show that $\sigma$ is surjective. Let $f\colon\mathrm{Spec}(R)\to\mathbb{F}_{2}$ be a continuous function. Then the set
$D(f)=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid f(\mathfrak{p})=1\}$
is a closed subset. Thus, D(f)D(f)is compact since $\mathrm{Spec}(R)$ is compact by Proposition 2.5. But $D(f)$ is also open, so it can be covered by open subsets of the form $D(a)$, where $a\in R$ according to the Zariski topology. Since $D(f)$ is compact, it can be covered by finitely many such, i.e.,
@@
$D(a)$ $=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a\not\in\mathfrak{p}\}$
$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a)\not\subseteq\mathfrak{p}\}$
$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a_{1},\ldots,a_{n})\not\subseteq\mathfrak{p}\}$
$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a_{i}\not\in\mathfrak{p}\textrm{ for some $i$}\}$
@@
so $D(f)=D(a)=D(\sigma(a))$. Since $f$ and $\sigma(a)$ take values in $\mathbb{F}_{2}$, we get that $f=\sigma(a)$. ∎
Notation 2.8.
@@
For every $x\in\mathbf{B}(X)$, we get from (2.5.1) for $R=\mathbf{B}(X)$ and $f=x$ an equality of sets
$\beta^{-1}(D(x))=\{p\in X\mid x(p)=1\}.$
Hence the preimage $\beta^{-1}(D(x))$ is open. Since the sets $\{D(x)\mid x\in\mathbf{B}(X)\}$ form a sub-basis of $\mathrm{Spec}(\mathbf{B}(X))$, we get that $\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X))$ is continuous.
@@
Since $X$ is compact and $\mathrm{Spec}(\mathbf{B}(X))$ is Hausdorff, it will be sufficient to show that $\beta$ is a bijection by [2, I §9.4, Corollary 2 on page 87], which states that a continuous map from a compact space to a Hausdorff space is a homeomorphism if it is a bijection. Let $x,y\in X$ be two distinct points. Since $X$ is totally separated, there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$. Let $f\colon X\to\mathbb{F}_{2}$ be the characteristic function of $U$. Thus $f(x)=1$ if $x\in U$ and $f(x)=0$ if $x\in X\setminus U=V$. Hence $f$ is continuous and an element in $\mathbf{B}(X)$. Clearly $f\in\beta(y)$, but $f\not\in\beta(x)$, so $\beta$ is injective.
For every ideal $I\triangleleft\mathbf{B}(X)$, let $Z(I)\subseteq X$ denote the closed subset
@@
so if this set is empty, then $f$ is the identically one function, and hence $I=(1)=\mathbf{B}(X)$ is not proper, a contradiction. Next consider the case when $I$ is finitely generated. Then $I$ is principal by part $(iv)$ of Proposition 2.3, so $Z(I)$ is non-empty by the above. Finally, consider the general case. Then $Z(I)$ is the intersection of sets of the form $Z(J)$ where $J$ is a finitely generated ideal of $I$. Since the latter collection of sets is closed under finite intersections, and each member is non-empty by the above, we get $Z(I)$ is also non-empty, since $X$ is compact.
Now let $\mathfrak{m}\triangleleft\mathbf{B}(X)$ be a maximal ideal. By the claim in the previous paragraph, there is an $x\in Z(\mathfrak{m})$. Clearly $\beta(x)\supseteq\mathfrak{m}$, but $\mathfrak{m}$ is maximal, and hence $\beta(x)=\mathfrak{m}$. Therefore $\beta$ is surjective as well. ∎
Notation 2.10.
@@
which is well-defined as we saw in Example 2.2, and
@@
Let $R$ be a Boolean ring. Then the following are equivalent: